OpenStudy (anonymous):

find the equation of a line that has an x intercept of 5 and passes through the point (2, -1) then tell me how you did it.

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) m=1/3

OpenStudy (anonymous):

find y intercept now

OpenStudy (anonymous):

after you have the slope you can use y-y1=m(x-x1) and get your line equation

OpenStudy (anonymous):

how did you find slope with only one point?

OpenStudy (anonymous):

unless (5, 0) is a point?

OpenStudy (anonymous):

when x=0 u get yint hence 1/3=(yi+1)/(0-2)

OpenStudy (anonymous):

0, 5 is also a point

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

how did you get 1/3 then :P? What two points did you use?

OpenStudy (anonymous):

solve for yi u get -5/3 plug it into original equation y=1/3x-5/3

OpenStudy (anonymous):

i used (0+1)/(5-2)

OpenStudy (anonymous):

0.o?

OpenStudy (anonymous):

do u follow?

OpenStudy (anonymous):

oh I kind of see.

OpenStudy (anonymous):

how did you know that (5 -2) would work?

OpenStudy (anonymous):

umm (y2-y1)/(x2-x1) is equation correct?

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

we have two point in the form (x,y) (5,0) and (2,-1)

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

okay. now that I recognize (5,0) as a point this is easy.

OpenStudy (anonymous):

>.< thank you.

OpenStudy (anonymous):

it's not letting me give you a medal, one sec.

OpenStudy (anonymous):

lol